OpenStudy (anonymous):

Verify my answer for: Find all solutions to the equation. cos^2x+2cosx+1=0

OpenStudy (anonymous):

Factor the equation. cos^2x+cosx+cosx+1=0 cosx(cosx+1)+(cosx+1)=0 (cosx+1)^2=0 Solve the quadratic. cosx+1=0 cosx=-1 x=pi

OpenStudy (anonymous):

Is this correct?

OpenStudy (tkhunny):

Are we restricted to \(x\in[0,2\pi)\)

OpenStudy (anonymous):

Sorry, I was away from my computer for a moment. It didn't say. @tkhunny

OpenStudy (tkhunny):

Oh, then you will have to be more general. \(x = \pi\) is only one of infinitely many solutions. Can you name them all with one expression?

OpenStudy (anonymous):

"pi*(n*2pi), n= 1, 2, 3, ..."? @tkhunny

OpenStudy (tkhunny):

Pretty close. \(\pi + other\;stuff\).

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (tkhunny):

Usually, \(\pi + 2k\pi\) for k and integer. This gets all the positive and negative multiples of \(2\pi\) from \(\pi\). You've only the positive multiples.